# Conditional regression for the Nonlinear Single-Variable Model

### v1

@@ A.1 Proof of Proposition 6
    Proof  Define intervals $I_{k}:=(-\sqrt{2(k+1)}\sigma_{\zeta},\sqrt{2k}\sigma_{\zeta}]\cup[\sqrt{2k}\sigma_{\zeta},\sqrt{2(k+1)}\sigma_{\zeta})$ for $k=0,1,2,\dots$. We first note that, thanks to ($\bm{\zeta}$ subG), we have $\zeta_{i}\in\bigcup_{k\leqslant\tau\log n}I_{k}$ with probability higher than $1-2n^{-\tau}$, for every $i=1,\dots,n$. Conditioned on this event and on $Y_{i}\in T$, $\Pi_{\gamma}X_{i}\in f^{-1}(T+\bigcup_{k\leq\tau\log n}I_{k})$. Meanwhile, $\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B},\zeta\in I_{k}\right]\in[\min f^{-1}(T+I_{k}),\max f^{-1}(T+I_{k})]$. It follows from assumption ($\bm{\omega_{f}}$) that we have an absolute bound upon conditioning on $Y_{i}\in T,\mathcal{B},\zeta_{i}\in\bigcup_{k\leq\tau\log n}I_{k}$:
@@
    $\lesssim$ $C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta}+\sigma_{\zeta}\sum_{k>\tau\log n}\sqrt{k}e^{-k})\lesssim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})$
@@ A.2 Proof of Proposition 7 and Corollary 8
    $\mathbb{P}\left\{\left\|\gamma(\Pi_{\gamma}X_{i})-\mathbb{E}[\gamma(\Pi_{\gamma}X_{i})\mid Y\in T,\mathcal{B}]\right\|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}$
    $\left\|\operatorname{{Cov}}[\gamma(\Pi_{\gamma}X)|Y\in T,\mathcal{B}]\right\|\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})$
@@
    $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X_{i})\rangle-\mathbb{E}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle\mid Y\in T,\mathcal{B}]\right|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}$
    $\mathrm{Var}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle|Y\in T,\mathcal{B}]\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})$
@@ A.3 Proof of Proposition 17 and Corollary 9
    Suppose (X subG), (Y subG), ($\bm{\zeta}$ subG), ($\bm{\gamma_{1}}$), (LCV), and ($\bm{\omega_{f}}$) hold true. Let ${\mu_{l,h}^{b}}$ be the mean of the $h$-th slice and ${v_{l,h}^{b}}$ be the significant vector of the $h$-th slice. Then, for every $l$ such that $l\gtrsim C_{f}C_{Y}R_{0}/\sigma_{\gamma}$, $|{R_{l,h}^{b}}|\geqslant\max\{\sigma_{\zeta},\omega_{f}\}$ for all $h\in\mathcal{H}_{l}^{b}$, for every $\epsilon\in(0,1)$ and $\tau\geqslant 1$, we have:
@@
    $\mathbb{P}\left(\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\ \Big{|}\ {n_{l,h}^{b}}\right)\lesssim d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(l^{-2}+l^{-1}\epsilon)}\right)+n^{-\tau}$
@@
    $\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|\leqslant\frac{\left\|({\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|}{\lambda_{d-1}({\widehat{\Sigma}_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})}$ (6)
@@
    $\lesssim$ $\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|+R_{0}\sqrt{d}\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|$
@@
    $\bm{\mathfrak{S}}_{h}=\left\{\begin{split}&\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|<\frac{R_{0}\sqrt{d}}{2},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\\ &\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|<\epsilon\sigma_{\gamma}^{2},\left|\widehat{V}_{l,h}^{b}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2}\end{split}\right\}$
@@
    $+\exp\left(-\frac{c\sigma_{\gamma}^{2}\beta^{2}{n_{l,h}^{b}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{2}l^{-4}(\beta l^{2}+C_{f}^{2}C_{Y}^{2}R_{0}^{2}\sigma_{\gamma}^{-2})}\right)+d\exp\left(-\frac{c\sigma_{\gamma}^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}$
@@ A.4 Proof of Proposition 10
    $Q_{h}(t_{0}):=\min\left(\mathbb{P}(t\in(0,t_{0})|Y\in{R_{l,h}})\,,\,\mathbb{P}(t\in(t_{0},\mathrm{len}_{\gamma})|Y\in{R_{l,h}})\right)$
@@
    $\gtrsim$ $C(c_{1})\min_{h={h^{\prime}_{x}},{h^{\prime}_{x}}-1}d(x,h)\gtrsim C(c_{1})C_{f}^{\prime 2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})$
@@ A.5 Proof of Proposition 11
    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\sum_{k^{\prime}=1}^{|k|}|\Delta\gamma_{{h_{x}}+k^{\prime}}|\cong|k|\mathrm{len}_{\gamma}/l$
@@
    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim K_{0}\sqrt{\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}}\mathrm{reach}_{\gamma}\gtrsim\frac{\mathrm{len}_{\gamma}}{l}\frac{\mathrm{reach}_{\gamma}}{\sigma_{\gamma}}\gtrsim\mathrm{len}_{\gamma}/l$
@@
    $\text{For any }h\text{ such that }|h-{h_{x}}|\geqslant 2,\ \sqrt{{\mathrm{dist}}(x,h)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\mathrm{len}_{\gamma}/l$
@@
    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|$ (7)
@@
    $\bm{\mathfrak{S}}(\epsilon,\delta,\beta,u)=\left\{\begin{split}&\text{ For all}h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\sigma_{\gamma}\sqrt{\delta},\\ &\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2},\left|{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\frac{u^{2}C_{f}^{2}C_{Y}^{2}R_{0}^{2}}{l^{2}}\end{split}\right\}$
@@
    $+ld\exp\left(-\frac{c\sigma_{\gamma}^{4}C_{Y}^{2}u^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{3/2}(\sqrt{d}+uC_{f}C_{Y})}\right)+l\exp\left(-\frac{c{n_{l,h}^{b}}u^{4}}{1+u^{2}}\right)$
@@
    $+\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}\left(\sigma_{\gamma}^{2}\delta+R_{0}\sigma_{\gamma}\sqrt{d}\sqrt{\delta}\right)+\frac{u^{2}\mathrm{len}_{\gamma}^{2}}{l^{2}}\frac{R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}}+\beta\sigma_{\gamma}^{2}\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}^{2}}$
@@ A.6 Proof of Proposition 12
    $\operatorname{\mathrm{MSE}}_{(NCA)}:=\mathbb{E}\left[\left|f(\Pi_{\gamma}X)-\widetilde{f}(\langle{v_{l,h}^{b}},X\rangle)\right|^{2}\mathbbm{1}_{I^{(l,h)}}(\langle{v_{l,h}^{b}},X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]$
@@
    $\begin{split}\operatorname{\mathrm{MSE}}_{(NCA)}\leqslant&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\Pi_{\gamma}X-\langle X,{v_{l,h}^{b}}\rangle-c(h)\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \leqslant&C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\langle\gamma(\Pi_{\gamma}X)-\gamma(t_{1}),{v_{l,h}^{b}}\rangle-(\Pi_{\gamma}X-t_{1})\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ &+C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|(Z_{d-1},0)^{\intercal}M_{\gamma^{\prime}(t)}^{\intercal}({v_{l,h}^{b}}-\gamma^{\prime}(t))\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \lesssim&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\sigma_{\gamma}^{2(s\wedge 1)}\left\|\gamma^{\prime\prime}\right\|^{2(s\wedge 1)}C_{f}^{2(s\wedge 1)}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2(s\wedge 1)}\end{split}\,.$ (NL)
@@
    $\begin{split}\operatorname{\mathrm{MSE}}_{(\Psi)}&:=\mathbb{E}\left[\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|^{2}\right]\end{split}$ ($\Psi$)

### v2

@@ A.1 Proof of Proposition 6
    Proof  Define intervals $I_{k}:=(-\sqrt{2(k+1)}\sigma_{\zeta},\sqrt{2k}\sigma_{\zeta}]\cup[\sqrt{2k}\sigma_{\zeta},\sqrt{2(k+1)}\sigma_{\zeta})$ for $k=0,1,2,\dots$. We first note that, thanks to ($\bm{\zeta}\in\psi_{2}$), we have $\zeta_{i}\in\bigcup_{k\leqslant\tau\log n}I_{k}$ with probability higher than $1-2n^{-\tau}$, for every $i=1,\dots,n$. Conditioned on this event and on $Y_{i}\in T$, $\Pi_{\gamma}X_{i}\in f^{-1}(T+\bigcup_{k\leq\tau\log n}I_{k})$. Meanwhile, $\mathbb{E}\left[\Pi_{\gamma}X|Y\in T,\mathcal{B},\zeta\in I_{k}\right]\in[\min f^{-1}(T+I_{k}),\max f^{-1}(T+I_{k})]$. It follows from assumption ($\bm{\omega_{f}}$) that we have an absolute bound upon conditioning on $Y_{i}\in T,\mathcal{B},\zeta_{i}\in\bigcup_{k\leq\tau\log n}I_{k}$:
@@
    $\lesssim$ $C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta}+\sigma_{\zeta}\sum_{k>\tau\log n}\sqrt{k}e^{-k})\lesssim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\ \,.$
@@ A.2 Proof of Proposition 7 and Corollary 8
    $\mathbb{P}\left\{\left\|\gamma(\Pi_{\gamma}X_{i})-\mathbb{E}[\gamma(\Pi_{\gamma}X_{i})\mid Y\in T,\mathcal{B}]\right\|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}\ \,;$
    $\left\|\operatorname{{Cov}}[\gamma(\Pi_{\gamma}X)|Y\in T,\mathcal{B}]\right\|\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})\ \,,$
@@
    $\mathbb{P}\left\{\left|\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X_{i})\rangle-\mathbb{E}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle\mid Y\in T,\mathcal{B}]\right|\gtrsim C_{f}(|T|+\sqrt{\tau\log n}\sigma_{\zeta})\mid Y_{i}\in T,\mathcal{B}_{i}\right\}\leqslant 2n^{-\tau}\ \,;$
    $\mathrm{Var}[\langle{v_{l,h}^{b}},\gamma(\Pi_{\gamma}X)\rangle|Y\in T,\mathcal{B}]\lesssim C_{f}^{2}(|T|^{2}+\sigma_{\zeta}^{2})\ \,.$
@@ A.3 Proof of Proposition 17 and Corollary 9
[…]
@@
    $\mathbb{P}\left(\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|>\frac{\sigma_{\gamma}^{2}\epsilon}{R_{0}\sqrt{d}}\ \Big{|}\ {n_{l,h}^{b}}\right)\lesssim d\exp\left(-\frac{c\sigma_{\gamma}^{4}\epsilon^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{4}d(l^{-2}+l^{-1}\epsilon)}\right)+n^{-\tau}\ \,.$
@@
    $\left\|{\widehat{v}_{l,h}^{b}}-{v_{l,h}^{b}}\right\|\leqslant\frac{\left\|({\widehat{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|}{\lambda_{d-1}({\widehat{\Sigma}_{l,h}^{b}})-\lambda_{d}({\Sigma_{l,h}^{b}})}\ \,.$ (6)
@@
    $\lesssim$ $\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|+R_{0}\sqrt{d}\left|\langle{v_{l,h}^{b}},{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\rangle\right|\ \,.$
@@
    $\bm{\mathfrak{S}}_{h}=\left\{\begin{split}&\left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|<\frac{R_{0}\sqrt{d}}{2},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\\ &\left\|({\widetilde{\Sigma}_{l,h}^{b}}-{\Sigma_{l,h}^{b}}){v_{l,h}^{b}}\right\|<\epsilon\sigma_{\gamma}^{2},\left|\widehat{V}_{l,h}^{b}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2}\end{split}\right\}\ \,.$
@@
    $+\exp\left(-\frac{c\sigma_{\gamma}^{2}\beta^{2}{n_{l,h}^{b}}}{C_{f}^{2}C_{Y}^{2}R_{0}^{2}l^{-4}(\beta l^{2}+C_{f}^{2}C_{Y}^{2}R_{0}^{2}\sigma_{\gamma}^{-2})}\right)+d\exp\left(-\frac{c\sigma_{\gamma}^{4}{n_{l,h}^{b}}}{R_{0}^{4}d^{2}}\right)+n^{-\tau}\ \,.$
@@ A.4 Proof of Proposition 10
    $Q_{h}(t_{0}):=\min\left(\mathbb{P}(t\in(0,t_{0})|Y\in{R_{l,h}})\,,\,\mathbb{P}(t\in(t_{0},\mathrm{len}_{\gamma})|Y\in{R_{l,h}})\right)\ \,.$
@@
    $\gtrsim$ $C(c_{1})\min_{h={h^{\prime}_{x}},{h^{\prime}_{x}}-1}d(x,h)\gtrsim C(c_{1})C_{f}^{\prime 2}(|{R_{l,h}}|^{2}+\sigma_{\zeta}^{2})\ \,.$
@@ A.5 Proof of Proposition 11
    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\sum_{k^{\prime}=1}^{|k|}|\Delta\gamma_{{h_{x}}+k^{\prime}}|\cong|k|\mathrm{len}_{\gamma}/l\ \,.$
@@
    $\sqrt{{\mathrm{dist}}(x,{h_{x}}+k)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim K_{0}\sqrt{\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}}\mathrm{reach}_{\gamma}\gtrsim\frac{\mathrm{len}_{\gamma}}{l}\frac{\mathrm{reach}_{\gamma}}{\sigma_{\gamma}}\gtrsim\mathrm{len}_{\gamma}/l\ \,.$
@@
    $\text{For any }h\text{ such that }|h-{h_{x}}|\geqslant 2,\ \sqrt{{\mathrm{dist}}(x,h)}-\sqrt{{\mathrm{dist}}(x,{h_{x}})}\gtrsim\mathrm{len}_{\gamma}/l\ \,.$
@@
    $\left|{\widehat{\mathrm{dist}}}(x,h^{\prime})-{\mathrm{dist}}(x,h^{\prime})\right|+\left|{\widehat{\mathrm{dist}}}(x,{h_{x}})-{\mathrm{dist}}(x,{h_{x}})\right|<\left|{\mathrm{dist}}(x,h^{\prime})-{\mathrm{dist}}(x,{h_{x}})\right|\ \,.$ (7)
@@
    $\bm{\mathfrak{S}}(\epsilon,\delta,\beta,u)=\left\{\begin{split}&\text{ For all}h\in{\mathcal{H}_{l}^{b}},\left|\langle{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}},{v_{l,h}^{b}}\rangle\right|\lesssim\frac{\epsilon\sigma_{\gamma}^{2}}{R_{0}\sqrt{d}},\ \left\|{\widehat{\mu}_{l,h}^{b}}-{\mu_{l,h}^{b}}\right\|\lesssim\sigma_{\gamma}\sqrt{\delta},\\ &\left\|{v_{l,h}^{b}}-{\widehat{v}_{l,h}^{b}}\right\|\lesssim\epsilon,\left|{\lambda_{1}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{1}(\Sigma_{l,h}^{b})}\right|\lesssim\beta\sigma_{\gamma}^{2},\left|{\lambda_{d}(\widehat{\Sigma}_{l,h}^{b})}-{\lambda_{d}(\Sigma_{l,h}^{b})}\right|\lesssim\frac{u^{2}C_{f}^{2}C_{Y}^{2}R_{0}^{2}}{l^{2}}\end{split}\right\}\ \,.$
@@
    $+ld\exp\left(-\frac{c\sigma_{\gamma}^{4}C_{Y}^{2}u^{2}{n_{l,h}^{b}}(\tau\log n)^{-\frac{1}{2}}}{R_{0}^{4}d^{3/2}(\sqrt{d}+uC_{f}C_{Y})}\right)+l\exp\left(-\frac{c{n_{l,h}^{b}}u^{4}}{1+u^{2}}\right)\ \,.$
@@
    $+\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}}{{\lambda_{1}(\Sigma_{l,h}^{b})}}\left(\sigma_{\gamma}^{2}\delta+R_{0}\sigma_{\gamma}\sqrt{d}\sqrt{\delta}\right)+\frac{u^{2}\mathrm{len}_{\gamma}^{2}}{l^{2}}\frac{R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}}+\beta\sigma_{\gamma}^{2}\frac{{\lambda_{d}(\Sigma_{l,h}^{b})}R_{0}^{2}d}{{\lambda_{1}(\Sigma_{l,h}^{b})}^{2}}\ \,.$
@@ A.6 Proof of Proposition 12
    $\operatorname{\mathrm{MSE}}_{(NCA)}:=\mathbb{E}\left[\left|f(\Pi_{\gamma}X)-\widetilde{f}(\langle{v_{l,h}^{b}},X\rangle)\right|^{2}\mathbbm{1}_{I^{(l,h)}}(\langle{v_{l,h}^{b}},X\rangle)\ \big{|}\ X\in{S_{l,h}}\right]\ \,.$
@@
    $\begin{split}\operatorname{\mathrm{MSE}}_{(NCA)}\leqslant&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\Pi_{\gamma}X-\langle X,{v_{l,h}^{b}}\rangle-c(h)\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \leqslant&C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|\langle\gamma(\Pi_{\gamma}X)-\gamma(t_{1}),{v_{l,h}^{b}}\rangle-(\Pi_{\gamma}X-t_{1})\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ &+C(s)\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\mathbb{E}\left[\left|(Z_{d-1},0)^{\intercal}M_{\gamma^{\prime}(t)}^{\intercal}({v_{l,h}^{b}}-\gamma^{\prime}(t))\right|^{2(s\wedge 1)}\ \big{|}\ X\in{S_{l,h}}\right]\\ \lesssim&\left[f\right]_{\mathcal{C}^{s\wedge 1}}^{2}\sigma_{\gamma}^{2(s\wedge 1)}\left\|\gamma^{\prime\prime}\right\|^{2(s\wedge 1)}C_{f}^{2(s\wedge 1)}\max\left(\sigma_{\zeta},\omega_{f},\frac{\mathrm{len}_{\gamma}}{C_{f}^{\prime}l}\right)^{2(s\wedge 1)}\ \,.\end{split}$ (NL)
@@
    $\begin{split}\operatorname{\mathrm{MSE}}_{(\Psi)}&:=\mathbb{E}\left[\left|{\widehat{f}_{j|{v_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)-{\widehat{f}_{j|{\widehat{v}_{l,h_{x}}^{b}}}}(\langle{\widehat{v}_{l,h_{x}}^{b}},x\rangle)\right|^{2}\right]\end{split}\ \,.$ ($\Psi$)